OpenStudy (anonymous):

verify using sin and cos (tanθ+cotθ)tanθ=sec^2θ

OpenStudy (anonymous):

convert it into terms of sin and cos

OpenStudy (anonymous):

You could use \[ \tan(\theta)\cot(\theta) = 1 \]

OpenStudy (anonymous):

You take advantage of (tanx)^2 = (secx)^2 - 1. This is a derivation of a pythagorean identity of trig functions. As @wio said, tant * cott = 1. Thus the sum is (secx)^2 + 1 - 1 = (secx)^2

OpenStudy (anonymous):

ok, I understand. But is there a way to use only sin and cos?

OpenStudy (anonymous):

\[\tan \theta = \frac{ \sin \theta }{\cos \theta }, \cot \theta = \frac{\cos \theta}{\sin \theta}\] so \[(\tan \theta + \cot \theta)\tan \theta = \frac{ \sin ^{2}\theta }{ \cos ^{2} \theta} + 1\] which equals \[\frac{ \sin^{2}\theta + \cos^{2} \theta }{\cos^{2}\theta }\] by the Pythagoean identity, we know that the numerator equals 1, thus we get \[\frac{ 1 }{ \cos^{2}\theta } = \sec^{2} \theta\] by definition

OpenStudy (anonymous):

ok, I got it. thanks for both of your help!